Solve the following problems.show the equation used for the solution.
Social science. In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?

Answer by ptaylor(2198)

Let x=number of votes for the loser
Then 220+x=number of votes for the winner
Total votes cast=810
Now we know that:
Number of votes for loser plus number of votes for winner equals total votes cast. So our equation to solve is:
x+220+x=810 subtract 220 from both sides
x+220-220+x=810-220 collect like terms
2x=590 divide both sides by 2
x=295--------------------------number of votes for the loser
220+x=220+295=515------------------number of votes for the winner

CK
295+515=810
810=810
also 515 is 220 more than 295

Hope this helps---ptaylor